A research company would like to determine the effectiveness of a pain-relief ointment designed for lower back pain. Prior to testing the ointment, each of the patients fill out a survey to rate their current level of back pain on a scale from 1 to 10. Then, they apply the ointment and after 90 minutes of the application, they fill out the survey a second time. The data are show in the table. If a two-tailed related samples t statistic at the .05 level of significance is computed, what is the value of the variance of the difference (s2D)? Enter the numerical value below. Include one digit after the decimal point.

Before After
6 2
8 3
9 4
8 1
10 2
5 3
9 8
7 7
Answer #1
Before After Difference
6 2 4 16
8 3 5 25
9 4 5 25
8 1 7 49
10 2 8 64
5 3 2 4
9 8 1 1
7 7 0 0

∑d = 32

∑d² = 184

n = 8

Mean , x̅d = Ʃd/n = 32/8 = 4

Variance of the difference, d = (Ʃd² - (Ʃd)²/n)/(n-1) = (184-(32)²/8)/(8-1) = 8.0